# Measuring Resolved Star Formation Histories from High-Precision Color-Magnitude Diagrams with StarFormationHistories.jl

### v1

@@ Abstract
Understanding how and when galaxies formed stars over the history of the Universe is fundamental to the study of galaxy evolution. The star formation histories (SFHs) of galaxies in the local Universe can be measured with high precision using deep imaging with space telescopes. Such resolved SFHs are based on modelling the observed color-magnitude diagram (CMD) with stellar evolution models and rely on age-sensitive features like the main sequence turn-off to measure a galaxy’s star formation rate as a function of time. There are many other population-level parameters that factor into these measurements, such as the stellar initial mass function (IMF), binary fraction, and metallicity, to name a few. We present and release StarFormationHistories.jl, a modular, open-source Julia package for measuring resolved SFHs with a focus on model flexibility for these types of population parameters. The code can model unresolved photometric binaries and supports arbitrary IMFs. Random uncertainties in the SFH measurements can be quantified with Monte Carlo posterior sampling methods. We illustrate the performance of the package on JWST/NIRCAM data of the Local Group dwarf irregular galaxy WLM $\left(M_{v}\approx-14.2\right)$, which exhibits a complex, well-sampled CMD, and HST/ACS data of the ultra-faint Milky Way satellite dwarf galaxy Horologium I $\left(M_{v}\approx-3.7\right)$, which has a much simpler but sparser CMD.
@@ 2.1 CMD Discretization
The most common approach to manual binning of the CMD is to impose a regular grid along the two component axes and count the number of stars that fall into each bin, creating effectively a two-dimensional histogram. Such a 2-D discretization of the CMD is known as a Hess diagram. With the CMD binned into a discrete Hess diagram, Poisson statistics can be used to quantify goodness-of-fit between the observed data and SSP models discretized to the same 2-D grid. This is the method used by the match software (Dolphin, 2002, 2012, 2013), one of the earliest and most prolific programs for measuring resolved SFHs (interesting applications of match include McQuinn et al., 2010; Weisz et al., 2011, 2014a, 2014b; Skillman et al., 2017; Hargis et al., 2020; Savino et al., 2023; McQuinn et al., 2024). We likewise use Hess diagrams with regular grids as our means of CMD discretization, as use of a regular grid enables specialized algorithms for generating model templates for SSPs (described in detail in §2.2).
@@ 2.5 Fitting Model Parameters
This exponential transformation requires marginally more BFGS iterations to achieve equivalent convergence criteria compared to $r_{j}=a_{j}^{2}$, but this minor increase in cost is offset by the fact that the exponential transformation allows us to obtain random uncertainty estimates on the $r_{j}$ from a BFGS optimization for free. We will present a short derivation here, but refer the reader to Dovì et al. (1991) and Appendix A of Yuen (2010) for more information. Recall that the Poisson distribution for large expectation value $\lambda$ is very similar in shape to a Gaussian distribution with mean $\lambda$ and variance $\lambda^{2}$. For a Gaussian loglikelihood objective function, it can be shown that the Hessian matrix is equal to the inverse of the Gaussian covariance matrix. Therefore the Hessian is constant for all parameters; since the loglikelihood is quadratic in the fitting variables, its second derivatives are all constants. As such, if one can calculate the Hessian from the Gaussian objective one can obtain the covariance matrix of the parameters in the fit. A similar calculation can be applied to other objectives to obtain approximations of the covariance matrices of their parameters. The validity of these approximations depends on the shape of the objective function in question.
@@ 2.7 Hierarchical Models
match supports a non-parametric AMR where $\langle[\text{M}/\text{H}]\rangle(t)$ is essentially unconstrained. It also implements simpler parametric forms that can be constrained to be monotonically increasing towards the present day; this is the approach we choose to take here. Although detailed chemical enrichment models indicate some Local Group dwarfs may have accreted small amounts of pristine gas at late times (e.g., Kirby et al., 2013), the assumption that the AMR increases monotonically to the present-day is generally acceptable for resolved SFH studies (see, e.g., appendix B of Savino et al., 2023). At each time $t$, match uses only a single template with age $t$ and metallicity $\langle[\text{M}/\text{H}]\rangle(t)$ to represent the stars forming at that time; i.e., the only spread in metallicity at fixed $t$ is whatever spread was chosen by the user when generating the template, which is typically 0.1–0.2 dex (e.g., Weisz et al., 2011). Both types of models greatly reduce the parameter space of possible template combinations by restricting the allowed spread in metallicity at fixed time. The non-parametric form allows additional freedom for the AMR to respond to properties of the SFH; for example, if the best-fit SFH results in a SFR of 0 between times $t_{1}$ and $t_{2}$ then the change in metallicity between these times can be 0 as well.
@@
Figure 6: Results of applying our SFH fitting methodology to the synthetic stellar population shown in Figure 5. The best-fit cumulative SFH (left) and age-metallicity relation (right) are consistent with those used to construct the synthetic population. The $68\%$ credible intervals are shaded in both panels, but this region is too small to see in the right panel.
@@ 3 Application to Synthetic Data
As the SFRs in adjacent time bins can be highly correlated (Dolphin, 2002), it is common to examine the cumulative SFH, defined as the fraction of stellar mass formed at times earlier than $t$. Such a cumulative statistic has the effect of canceling out the correlations between adjacent time bins that can be significant when looking directly at SFRs. The left panel of Figure 6 shows our best-fit cumulative SFH for this synthetic population with the $68\%$ credible interval shaded. Our result is statistically consistent with the intrinsic SFH of the population. The right panel shows our best-fit AMR for the synthetic population, which is also in very good agreement with the intrinsic AMR. The $68\%$ credible interval is also shaded in the right panel, but it is too narrow to see.
@@
In summary, we are able to accurately recover the input SFH for our synthetic population. This is most obvious in the best-fit cumulative SFH and AMR shown in Figure 6, as these are less affected by specifics of the analysis (e.g., the time resolution of the adopted isochrone grid) than the SFRs shown in Figure 7. Even so, the best-fit SFRs show good statistical agreement with the input value of $\sim 0.7\times 10^{-3}$ M yr-1, giving us confidence to apply our methodology on real data.
@@
Figure 9: Left: Cumulative SFHs of WLM measured from the JWST/NIRCAM photometric catalogs of Weisz et al. (2024) with our methodology (blue) and match (orange, Cohen et al., in preparation). The $68\%$ credible interval for our result, considering only random uncertainty, is shaded grey. While there are minor differences, the two results are generally in good agreement. The shaded red region shows the $68\%$ confidence interval for the midpoint of reionization (Planck Collaboration et al., 2020). Right: Solutions for the age-metallicity relation found with our methodology (blue) and match (orange, Cohen et al., in preparation). The slopes are in excellent agreement, with only a minor $\sim 0.1$ dex constant offset. This is likely due to differences in how we model the fixed-age metallicity distribution.
@@ 4 WLM
Since McQuinn et al. (2024) used pre-release photometric catalogs that are not publicly available and focused their analysis on the region of WLM where the HST/ACS and JWST/NIRCAM imaging overlap, we do not compare directly to their published results. Rather, we compare against new match results that use the public photometric catalogs and ASTs released as part of the JWST Resolved Stellar Populations Early Release Science Program (Weisz et al., 2024) for the full JWST/NIRCAM field of view (Cohen et al., in preparation). The modelling assumptions and match parameters used to make these measurements are the same as were used in McQuinn et al. (2024), with the only difference being the input photometric catalogs and ASTs. These results are qualitatively consistent with those published in McQuinn et al. (2024), as differences between the pre-release catalogs used by McQuinn et al. (2024) and the public catalogs are minimal. We use the same public photometric catalogs (Weisz et al., 2024) and both analyses apply the star-galaxy separation criteria from Warfield et al. (2023). As McQuinn et al. (2024) demonstrated that the resolved SFHs derived from the HST/ACS data and the JWST/NIRCAM data are consistent, we consider only the JWST/NIRCAM data here. We apply our methodology to HST/ACS data of the Horologium I UFD in §5 so that we have an example of our methodology applied to both observatories.
@@
This consistency is reflected in the measured cumulative SFH shown in the left panel of Figure 9. While there are minor differences at intermediate ages (e.g., 3–5 Gyr), our measurement is broadly consistent with that from match (Cohen et al., in preparation). In particular, agreement is excellent within the last 3 Gyr, and we also replicate the quiescent period found by match following reionization ($z_{re}=7.67\pm 0.73$, $t_{re}=13.11\pm 0.09$ Gyr, Planck Collaboration et al., 2020). A useful comparison is figure 9 of McQuinn et al. (2024) which shows cumulative SFHs for WLM fit with different data (JWST/NIRCAM, HST/ACS, and a joint fit) and different stellar models – the variance between these results is significantly greater than the discrepancies we see between our result and the match result in Figure 9.
@@
The right panel of Figure 9 compares the match AMR to our best-fit linear AMR of $\langle[\text{M}/\text{H}]\rangle(t)=0.07\left(13.7-t\right)-1.44$ where $t$ is the lookback time in units of Gyr. This slope is consistent with the match result while the intercept is $\sim 0.1$ dex larger, likely due differences in how we model the fixed-age metallicity distribution. In any case, 0.1 dex is quite minor as metallicity inference from broadband photometry (particularly with only one color) is typically imprecise. Overall we find that our solution for WLM is in good agreement with the results from match (McQuinn et al., 2024, Cohen et al., in preparation), despite the differences in our methodologies.
@@ 5.1 Technical Considerations
All previously studied UFDs host ancient stellar populations of ages $\geq 10$ Gyr (Brown et al., 2014; Weisz et al., 2014a, b; Sacchi et al., 2021; Savino et al., 2023), which, in combination with their intrinsically low luminosities ($M_{V}>-7$ mag), makes them uniquely challenging to model. As previously mentioned, the SFRs of old stellar populations can often have strong covariances due to how few age-sensitive CMD features there are in these old stellar populations. The primary CMD feature that provides age sensitivity for old, metal-poor populations is the MSTO – the ability to fit the resolved SFH of a UFD is therefore highly dependent on the presence of a well-resolved and reasonably-sampled MSTO.
@@ 6 Discussion
We have so far neglected discussing systematic uncertainties because we have tried to make the same assumptions in our analyses as the literature work that we have been comparing to. However, for future, original analyses we will need to address systematics. For resolved SFHs, the dominant systematic uncertainties come from the stellar evolutionary tracks and bolometric corrections used to model the observations (Gallart et al., 2005; Dolphin, 2016), which is why special attention should be paid to the stellar libraries used for these analyses. It is common to constrain these systematic uncertainties by measuring the resolved SFH from the same observational data with multiple stellar libraries (e.g., Skillman et al., 2017; Sacchi et al., 2021), the idea being that systematic differences in the underlying stellar tracks and BCs will propagate to systematic differences in the resolved SFH measurement. It is easy to use isochrones from different stellar libraries with our code, making this form of systematic uncertainty estimation simple.

### v2

@@ Abstract
Understanding how and when galaxies formed stars over the history of the Universe is fundamental to the study of galaxy evolution. The star formation histories (SFHs) of galaxies in the local Universe can be measured with high precision using deep imaging with space telescopes. Such resolved SFHs are based on modelling the observed color-magnitude diagram (CMD) with stellar evolution models and rely on age-sensitive features like the main sequence turn-off to measure a galaxy’s star formation rate as a function of time. There are many other population-level parameters that factor into these measurements, such as the stellar initial mass function (IMF), binary fraction, and metallicity, to name a few. We present and release StarFormationHistories.jl, a modular, open-source Julia package for measuring resolved SFHs with a focus on model flexibility for these types of population parameters. The code can model unresolved photometric binaries and supports arbitrary IMFs. Random uncertainties in the SFH measurements can be quantified with Monte Carlo posterior sampling methods. We illustrate the performance of the package on JWST/NIRCAM data of the Local Group dwarf irregular galaxy WLM $\left(M_{v}\approx-14.2\ \text{mag}\right)$, which exhibits a complex, well-sampled CMD, and HST/ACS data of the ultra-faint Milky Way satellite dwarf galaxy Horologium I $\left(M_{v}\approx-3.7\ \text{mag}\right)$, which has a much simpler but sparser CMD.
@@ 2.1 CMD Discretization
The most common approach to manual binning of the CMD is to impose a regular grid along the two component axes and count the number of stars that fall into each bin, creating effectively a two-dimensional histogram. Such a 2-D discretization of the CMD is known as a Hess diagram. With the CMD binned into a discrete Hess diagram, Poisson statistics can be used to quantify goodness-of-fit between the observed data and SSP models discretized to the same 2-D grid. This is the method used by the match software (Dolphin, 2002, 2012, 2013), one of the earliest and most prolific programs for measuring resolved SFHs (interesting applications of match include McQuinn et al., 2010a; Weisz et al., 2011, 2014a, 2014b; Skillman et al., 2017; Hargis et al., 2020; Savino et al., 2023; McQuinn et al., 2024). We likewise use Hess diagrams with regular grids as our means of CMD discretization, as use of a regular grid enables specialized algorithms for generating model templates for SSPs (described in detail in §2.2).
@@ 2.5 Fitting Model Parameters
This exponential transformation requires marginally more BFGS iterations to achieve equivalent convergence criteria compared to $r_{j}=a_{j}^{2}$, but this minor increase in cost is offset by the fact that the exponential transformation allows us to obtain random uncertainty estimates on the $r_{j}$ from a BFGS optimization for free. We will present a short derivation here, but refer the reader to Dovì et al. (1991) and Appendix A of Yuen (2010) for more information. Note that the Poisson distribution for large expectation value $\lambda$ is very similar in shape to a Gaussian distribution with mean $\lambda$ and variance $\lambda^{2}$ (see, for example, equation 11.11 of Taylor, 1997). For a Gaussian loglikelihood objective function, it can be shown that the Hessian matrix is equal to the inverse of the Gaussian covariance matrix. Therefore the Hessian is constant for all parameters; since the loglikelihood is quadratic in the fitting variables, its second derivatives are all constants. As such, if one can calculate the Hessian from the Gaussian objective one can obtain the covariance matrix of the parameters in the fit. A similar calculation can be applied to other objectives to obtain approximations of the covariance matrices of their parameters. The validity of these approximations depends on the shape of the objective function in question.
@@ 2.7 Hierarchical Models
match supports a non-parametric AMR where $\langle[\text{M}/\text{H}]\rangle(t)$ is essentially unconstrained. It also implements simpler parametric forms that can be constrained to be monotonically increasing towards the present day; this is the approach we choose to take here. Although detailed chemical enrichment models indicate some Local Group dwarfs may have accreted small amounts of pristine gas at late times (e.g., Kirby et al., 2013), the assumption that the AMR increases monotonically to the present-day is generally acceptable for resolved SFH studies (see, e.g., appendix B of Savino et al., 2023). At each time $t$, match uses only a single template with age $t$ and metallicity $\langle[\text{M}/\text{H}]\rangle(t)$ to represent the stars forming at that time; i.e., the only spread in metallicity at fixed $t$ is whatever spread was chosen by the user when generating the template, which is typically 0.1–0.2 dex (e.g., Weisz et al., 2011), such that the number of free metallicity-related parameters is equal to the number of age bins. Both types of models greatly reduce the parameter space of possible template combinations by restricting the allowed spread in metallicity at fixed time. The non-parametric form allows additional freedom for the AMR to respond to properties of the SFH; for example, if the best-fit SFH results in a SFR of 0 between times $t_{1}$ and $t_{2}$ then the change in metallicity between these times can be 0 as well.
@@
Figure 6: Results of applying our SFH fitting methodology to the synthetic stellar population shown in Figure 5. The best-fit cumulative SFH (left) and age-metallicity relation (right) are consistent with those used to construct the synthetic population. The $68\%$ credible intervals, considering only random uncertainties, are shaded in both panels, but this region is too small to see in the right panel.
@@ 3 Application to Synthetic Data
As the SFRs in adjacent time bins can be highly correlated (Dolphin, 2002; McQuinn et al., 2010b), it is common to examine the cumulative SFH, defined as the fraction of stellar mass formed at times earlier than $t$. Such a cumulative statistic has the effect of canceling out the correlations between adjacent time bins that can be significant when looking directly at SFRs. The left panel of Figure 6 shows our best-fit cumulative SFH for this synthetic population with the $68\%$ credible interval shaded. Our result is statistically consistent with the intrinsic SFH of the population. The right panel shows our best-fit AMR for the synthetic population, which is also in very good agreement with the intrinsic AMR. The $68\%$ credible interval is also shaded in the right panel, but it is too narrow to see.
@@
In summary, we are able to accurately recover the input SFH for our synthetic population. This is most obvious in the best-fit cumulative SFH and AMR shown in Figure 6, as these are less affected by specifics of the analysis (e.g., the time resolution of the adopted isochrone grid) than the SFRs shown in Figure 7. Even so, the best-fit SFRs show good statistical agreement with the input value of $\sim 0.7\times 10^{-3}$ M yr-1, giving us confidence to apply our methodology on real data. We include synthetic data experiments in Appendix B that illustrate the performance of the SFH measurements on more distant galaxies without resolved MSTOs. We additionally examine the impact of observational systematics (particularly uncertainties on distances, foreground extinctions, and stellar binary fractions) on our ability to recover cumulative SFHs and AMRs in Appendix C.
@@
Figure 9: Left: Cumulative SFHs of WLM measured from the JWST/NIRCAM photometric catalogs of Weisz et al. (2024) with our methodology (blue) and match (orange, Cohen et al., 2025). The $68\%$ credible interval for our result, considering only random uncertainty, is shaded grey. While there are minor differences, the two results are generally in good agreement. The shaded red region shows the $68\%$ confidence interval for the midpoint of reionization (Planck Collaboration et al., 2020). Right: Solutions for the age-metallicity relation found with our methodology (blue) and match (orange, Cohen et al., 2025). The slopes are in excellent agreement, with only a minor $\sim 0.1$ dex constant offset. This is likely due to differences in how we model the fixed-age metallicity distribution.
@@ 4 WLM
Since McQuinn et al. (2024) used pre-release photometric catalogs that are not publicly available and focused their analysis on the region of WLM where the HST/ACS and JWST/NIRCAM imaging overlap, we do not compare directly to their published results. Rather, we compare against new match results that use the public photometric catalogs and ASTs released as part of the JWST Resolved Stellar Populations Early Release Science Program (Weisz et al., 2024) for the full JWST/NIRCAM field of view (Cohen et al., 2025). The modelling assumptions and match parameters used to make these measurements are the same as were used in McQuinn et al. (2024), with the only difference being the input photometric catalogs and ASTs. These results are qualitatively consistent with those published in McQuinn et al. (2024), as differences between the pre-release catalogs used by McQuinn et al. (2024) and the public catalogs are minimal. We use the same public photometric catalogs (Weisz et al., 2024) and both analyses apply the star-galaxy separation criteria from Warfield et al. (2023). As McQuinn et al. (2024) demonstrated that the resolved SFHs derived from the HST/ACS data and the JWST/NIRCAM data are consistent, we consider only the JWST/NIRCAM data here. We apply our methodology to HST/ACS data of the Horologium I UFD in §5 so that we have an example of our methodology applied to both observatories.
@@
This consistency is reflected in the measured cumulative SFH shown in the left panel of Figure 9. While there are minor differences at intermediate ages (e.g., 3–5 Gyr), our measurement is broadly consistent with that from match (Cohen et al., 2025). In particular, agreement is excellent within the last 3 Gyr, and we also replicate the quiescent period found by match following reionization ($z_{re}=7.67\pm 0.73$, $t_{re}=13.11\pm 0.09$ Gyr, Planck Collaboration et al., 2020). A useful comparison is figure 9 of McQuinn et al. (2024) which shows cumulative SFHs for WLM fit with different data (JWST/NIRCAM, HST/ACS, and a joint fit) and different stellar models – the variance between these results is significantly greater than the discrepancies we see between our result and the match result in Figure 9.
@@
The right panel of Figure 9 compares the match AMR to our best-fit linear AMR of $\langle[\text{M}/\text{H}]\rangle(t)=0.07\left(13.7-t\right)-1.44$ where $t$ is the lookback time in units of Gyr. This slope is consistent with the match result while the intercept is $\sim 0.1$ dex larger, likely due differences in how we model the fixed-age metallicity distribution. In any case, 0.1 dex is quite minor as metallicity inference from broadband photometry (particularly with only one color) is typically imprecise.
While we recover a cumulative SFH for WLM that matches well with the qualitative features of the match result, it is instructive to consider what quantiative differences may exist more closely. For this purpose, we present the raw SFH of WLM in Figure 10 to supplement the cumulative SFH shown in Figure 9. The measurement from match is also shown for comparison. We note that the match solution for times more recent than 1 Gyr has a time resolution of $\Delta\log{\left(\text{age}\right)}=0.1$ dex and changes to $\Delta\log{\left(\text{age}\right)}=0.05$ dex for earlier times, while our solution uses $\Delta\log{\left(\text{age}\right)}=0.05$ dex throughout. The match uncertainties (per age bin) are therefore smaller than ours for times more recent than 1 Gyr. The effect of reducing the time resolution on the SFR measurement uncertaintes is shown in Figure 7.
For times more recent than 1 Gyr $\left(\log{\left(\text{age}\right)}<9\right)$ we see very good agreement between our measurement and that from match which we attribute to the well-sampled upper MS allowing for a robust measurement. Agreement is good from 1 – 4 Gyr ago $\left(9\leq\log{\left(\text{age}\right)}\leq 9.6\right)$, with some SFR transposition due to temporal covariance. As previously mentioned, SFRs in adjacent time bins can be highly correlated (Dolphin, 2002; McQuinn et al., 2010b), an effect which is mitigated in a cumulative statistic like the cumulative SFH. In the raw SFRs we are now considering, these covariances do not cancel out, but cannot be easily visualized either – the error bars are typically standard deviations and so do not convey these covariances. The SFRs in the range of 1 – 4 Gyr ago exhibit the effects of this covariance, where under (over) estimates of the SFR relative to match are generally offset by over (under) estimates in an adjacent bin.
The only time period where we see divergence from match not explained by inter-bin covariance is in the period from 4 – 6 Gyr ago $\left(9.6\leq\log{\left(\text{age}\right)}\leq 9.8\right)$. Additionally, both solutions show a large degree of time variability in the SFR during this time period. We believe these are related – populations with ages 4 – 6 Gyr under our fiducial AMR model have significant red clump features, and while the red clump of WLM is very well-populated, it is difficult to model the red clump to the precision of space-based observatories like HST/ACS (see, e.g., Figure 4 of Albers et al., 2019) and JWST/NIRCam (see, e.g., Figure 6a of McQuinn et al., 2024, and our Figure 8). As such, we hypothesize that the discrepancies in the SFRs 4 – 6 Gyr ago are driven by difficulties modelling the red clump features of populations with these ages. In such cases where particular CMD features present modelling challenges, a CMD discretization technique that can separate the CMD by feature (such as that used by IAC-pop; Aparicio & Hidalgo, 2009) would be useful. One could imagine a scheme by which the MS and subgiant branch may be given increased statistical weight in the fit relative to evolved star features (e.g., the red clump and horizontal branch) could mitigate some of these challenges. We leave exploration of this to future work.
Again, the cumulative SFH we measure for WLM is in good agreement with the results from match (McQuinn et al., 2024; Cohen et al., 2025), illustrating all the same qualitative features. Quantitatively, the raw SFRs we measure are also in good agreement with match across most of cosmic time, with the exception of a small period of divergence 4 – 6 Gyr ago which we attribute to challenges in modelling the red clump of these populations at the level of precision achieved by our JWST/NIRCam photometry. These results give us confidence in our methodology and form an important point of reference for comparing results obtained with our methodology to those measured with match, which has been used to make many key measurements in the field.
Figure 10: The global SFH of WLM measured from the JWST/NIRCam data with our method (blue bars, black error bars) compared to the measurement on the same data with match (orange bars, red error bars; Cohen et al., 2025). Error bars shown are $68\%$ credible intervals considering only random (statistical) uncertainty. Agreement is excellent from the present-day to 1 Gyr ago ($x=9$) and good from 1 Gyr ago to about 4 Gyr ago ($x=9.6$). We see large covariances for time bins from 4 – 6 Gyr ago ($x=$9.6–9.8) that result in worse agreement with the match result. We attribute these differences to the significant red clump features present in stellar populations of these ages under our fiducial AMR – the red clump is known to be difficult to model to the level of precision achieved by our data.
@@ 5.1 Technical Considerations
All previously studied UFDs host ancient stellar populations of ages $\geq 10$ Gyr (Brown et al., 2014; Weisz et al., 2014a, b; Sacchi et al., 2021; Savino et al., 2023), which, in combination with their intrinsically low luminosities ($M_{V}>-7$ mag), makes them uniquely challenging to model. While their stellar populations are ancient, UFDs require a spread in stellar ages to explain their CMDs as measured with high-precision space-based observatories; i.e., they are not mono-age populations. Additionally, they have significant intrinsic metallicity spreads of 0.2–0.7 dex (e.g., Kirby et al., 2011; Fu et al., 2023). Given these factors, the CMDs of these objects are best analyzed with a SFH fitting code rather than simpler isochrone fitting techniques often applied to globular clusters (e.g., VandenBerg et al., 2013, 2022). As previously mentioned, the SFRs of old stellar populations can often have strong covariances due to how few age-sensitive CMD features there are in these old stellar populations. The primary CMD feature that provides age sensitivity for old, metal-poor populations is the MSTO – the ability to fit the resolved SFH of a UFD is therefore highly dependent on the presence of a well-resolved and reasonably-sampled MSTO.
@@ 6 Discussion
We have so far neglected discussing systematic uncertainties because we have tried to make the same assumptions in our analyses as the literature work that we have been comparing to. However, for future, original analyses we will need to address systematics. For resolved SFHs, the dominant systematic uncertainties come from the stellar evolutionary tracks and bolometric corrections used to model the observations (Gallart et al., 2005; Dolphin, 2012, 2016), which is why special attention should be paid to the stellar libraries used for these analyses. It is common to constrain these systematic uncertainties by measuring the resolved SFH from the same observational data with multiple stellar libraries (e.g., Skillman et al., 2017; Sacchi et al., 2021), the idea being that systematic differences in the underlying stellar tracks and BCs will propagate to systematic differences in the resolved SFH measurement. It is easy to use isochrones from different stellar libraries with our code, making this form of systematic uncertainty estimation simple.